# Bias in Language Models: Beyond Trick Tests and Toward RUTEd Evaluation

### v1

@@ 3.2.1 Bedtime Stories
We include “Once upon a time” after initial trials showed it to be effective in inducing the model to immediately generate a story rather than generate text about bedtime stories more generally or conversational responses. We used a maximum length of 1000 tokens, which would be around 60 to 90 seconds spoken aloud at a slow-to-medium pace.
3.2.2 User Personas
An increasingly popular and influential use case for LLMs has been generating synthetic data that approximates human behavior, such as in psychology research (Crockett & Messeri, 2023; Dillion et al., 2023; Harding et al., 2023). In human-computer interaction, researchers have been exploring the interaction between LLMs as a data source, including for collective behavior such as creating a Reddit-like platform to simulate different community guidelines (Hämäläinen et al., 2023; Park et al., 2022, 2023). User research typically involves user personas that allow consideration of different user groups. They are short character sketches of fictional people that can include information such as a name, profile picture, basic demographics (age, gender, education level, etc.), hobbies, personal interests, accessibility needs, intended use of the product, and beliefs. Bias in this context could impact whose needs are considered as human behavior is understood and products are designed, built, and deployed. This risks systematically leading to real-world gaps in who is best served by new products and technology. This risks similar harms to those created by the longstanding gender and race biases in medical research and product development that have contributed to health disparities (Hoffman et al., 2016; Murthy et al., 2004).
To generate user personas amenable to evaluation of gender-occupation bias, we use the following prompt to elicit user personas for people of different occupations using an unspecified LLM-based product:
@@ 3.2.2 User Personas
For this task, we use a maximum length of 150 tokens to reflect the typical length in practice. This gave us a fair amount of detail, though in many cases, some content was cut off at the end. Because we knew from the task that the text should be focused on a single individual, we did not need completed text to extract the required information to calculate bias metrics.
3.2.3 ESL learning exercises
One of the most prevalent and widely discussed LLM uses is education; for example, the popular online learning platform Khan Academy has heavily invested in Khanmigo, a GPT-4-powered learning assistant (OpenAI, 2023). One important use case amenable to study is the generation of instructional content such as explanations, examples, and test questions. We prompt the model to generate sentences that could be used in instruction to demonstrate the use of pronouns to an English language learner (English as a Second Language; ESL). In this case, a significant gender imbalance could have similar harms as those to bedtime stories for children or for adults who are getting to know the culture and social norms of English-speaking regions, as well as harm through language learning itself as students learn different aspects of the language at different rates and learn to favor certain words in certain contexts. Specifically, we input the prompt:
@@ 3.2.3 ESL learning exercises
For this task, we use a maximum length of 100 tokens, corresponding to typical usage. This often resulted in a mid-sentence cutoff for the last sentence, but again, the text was sufficient to determine the gendered pronouns being used. This also typically produced several short, complete sentences sufficient for use in this context if the last incomplete sentence is dropped.
@@ 3.3 Metrics
Across each of these tasks, we apply the same metrics. For the decontextualized tasks, we are able to extract the probability of response directly. For instance, for the WinoBias-based task, for each sentence $s$, we extract the probability of the correct referent if the sentence is gender-occupation stereotypical, $p_{s}^{\text{s}}$, or the probability of the correct pronoun referent if the sentence is gender-occupation anti-stereotypical, $p_{s}^{\text{a}}$. Similarly, for the simple sentence-completion task, as discussed, for each occupation, $o$, we can directly extract the normalized probability that the next word in the sentence was a “male” word, $p_{o}^{m}$, or a “female” word, $p_{o}^{f}$.
@@
For the content generation tasks, we have to estimate the probabilities rather than extract them directly. For the Llama-2 models, we generate $n=30$ completions per task and occupation; for the PaLM models we generate $n=64$. Then, for each occupation, $o$, we calculate the proportion of generations that were about males, $\hat{p}_{o}^{m}$, and females, $\hat{p}_{p}^{f}$. We determine whether a generation is about a male or female by counting the number of male/female pronouns in the text. Replicates that contain no such pronouns are dropped. Those for which greater than half of the pronouns refer to males are categorized as “male” replicates; the others are categorized as “female.” Similarly, because each of the occupations we used has an associated gender-majority, we also calculate the proportion of replicates that were gender-stereotypical for the given occupation, $\hat{p}_{o}^{s}$ and gender anti-stereotypical, $\hat{p}_{o}^{a}$.
Going forward in defining the metrics, we drop the hat notation. For the tasks for which we can directly extract the probabilities, we plug in $p$, and for those for which we must estimate a proportion, we plug in $\hat{p}$.
3.3.1 Stereotype
We define our stereotype metric as follows:
$m^{\text{stereotype}}=\frac{1}{I}\sum_{i}p_{i}^{s}-p_{i}^{a}.$
To create an trick test evaluation evaluation, we apply this metric to the WinoBias task, and the sum ranges over sentences in the benchmark. Here, $I$ represents the number of sentences in the corpus. We also apply this metric to each of the RUTEd tasks, and $i$ ranges over occupations. For the RUTEd evaluations, $I$ represents the number of occupations considered (40).
For the WinoBias evaluation, a positive value for this metric means the normalized probability of the correct answer is higher when the sentence is consistent with existing gender stereotypes; a negative value for this metric means that the probability of the correct answer higher when the sentence is inconsistent with existing gender stereotypes. Similarly, for the content generation tasks, a higher value means that the gender distribution in occupations was more consistent with stereotypes; a negative value indicates the gender distribution of generated text was inconsistent with stereotypes.
3.3.2 Neutrality
We define
$m^{\text{neutrality}}=\frac{1}{O}\sum_{o}\left|p^{m}_{o}-p^{f}_{o}\right|.$
@@ 3.3.2 Neutrality
This metric is that suggested in the BigBench Gender Sensitivity-English task. Here, $O$ represents the number of occupations used in the task. Essentially, this measures a distance from parity. When applying this metric to the decontextualized sentence completion task, this metric is zero if the male words and female words have equal probability of coming next in the sentence. When applying this metric to the RUTEd text-generation tasks, this metric is zero if male and female replicates are equally likely to be generated.
3.3.3 Skew
We compute a similar skew metric. Instead of taking the absolute difference from parity, this metric is the average tendency of the model to return male words/replicates instead of female words/replicates.
$m^{\text{skew}}=\frac{1}{O}\sum_{o}\left(p^{\text{m}}_{o}-p^{\text{f}}_{o}\right).$
@@ 3.3.3 Skew
If male words/replicates have a higher probability systematically across all considered occupations, this metric is positive. Conversely, if female words/replicates have a systematically higher probability, the metric is negative.
@@ 3.4 Statistical Uncertainty
Because the probabilities in the decontextualized evaluations are directly observed, there is no statistical uncertainty, as we consider the set of prompts to be fixed. For the RUTEd evaluations, we estimate the sampling variance of each estimated probability. To properly propagate this uncertainty, we first note that one component used in all of the metrics is $d_{o}=\hat{p}_{o}^{m}-\hat{p}_{o}^{f}=\hat{p}_{o}^{m}-(1-\hat{p}_{o}^{m})$. For the RUTEd tasks, using a simple plug-in estimator of the sampling variance of $\hat{p}_{o}^{m}$, we get $d_{o}=2\hat{p}_{o}^{m}-1$ with sampling variance $\hat{\sigma}_{o}^{2}=4\frac{\hat{p}_{o}^{m}(1-\hat{p}_{o}^{m})}{n}$.
3.4.1 Neutrality
A rule of thumb for assuming approximate normality of $\hat{p}$ for a binomial distribution is that if you have at least ten positive and negative examples, this estimate is approximately correct (Peizer & Pratt, 1968). In our case, we largely meet or surpass this standard, so for mathematical convenience, we proceed under the assumption of normality—specifically, that $d_{o}\sim N(\mu_{o}=2\hat{p}_{o}^{m}-1,\hat{\sigma}^{2}_{o})$. This implies that $|d_{o}|$ has a folded normal distribution with parameters $\mu_{Y,o}=\hat{\sigma}_{o}{\sqrt{\tfrac{2}{\pi}}}\,e^{(-\mu_{o}^{2}/2\ \hat{\sigma_{o}}^{2})}+\mu_{o}\left(1-2\,\Phi(-{\tfrac{\mu_{o}}{\hat{\sigma}_{o}}})\right)$ and variance $\sigma_{Y,o}=\mu_{o}^{2}+\ \hat{\sigma}^{2}_{o}-\mu_{Y,o}^{2}$. This implies that the sampling variance of $m_{r}^{\text{neutrality}}$ (where $r$ denotes a RUTEd evaluation) is given by $\frac{1}{O^{2}}\sum_{o}\sigma_{Y,o}$.
3.4.2 Skew and Stereotype
Calculating the sampling variance of $m_{g}^{\text{skew}}$ and $m_{g}^{\text{stereotype}}$ is straightforward given $\hat{\sigma}_{o}^{2}$, as we are just averaging across independent approximate normal distributions. Therefore, in both cases, the sampling variance of the estimator is $\frac{1}{O^{2}}\sum\hat{\sigma_{o}}^{2}$.
@@ 4 Results
We generated content and calculated metrics for models from two different families: Llama-2 and Flan-PaLM. In both cases we used the chat/instruction tuned versions. For Llama 2’s chat-tuned models we evaluate 7, 13, and 70 billion parameters (Touvron et al., 2023) models, and for Flan-PaLM, we evaluate the XS, S, M, and L sizes (Chung et al., 2022). The results are shown in Figure 1. Columns of the grid correspond to metric types (i.e., neutrality, skew, and stereotype), and rows correspond to contexts (i.e., decontextualized, bedtime stories, personas, and ESL).
Figure 1: Bias evaluations for several sizes of Llama-2 (blue) and PaLM (orange) models.
Overall, the results confirm that the bias metrics that we computed are not measuring the same construct across different evaluation tasks. For example, among the Llama 2 models, for each metric type, we see that if we had used the decontextualized evaluation to select which of the model sizes is the least biased, this only would have corresponded with the optimal model in terms of a RUTEd evaluation as often as chance. Specifically, across all three decontextualized evaluations, the 13B model exhibited the least bias; across the neutrality metric, the 13B model only exhibited the least bias in one context, ESL; across the skew metric, the 13B model exhibited the least bias in one context, bedtime stories; and across the stereotype metric, the 13B model did not exhibit the least bias in any RUTEd context.
Patterns are qualitatively similar in the Flan PaLM2 models. Notably, there is little concordance across RUTEd evaluations as well, suggesting little generalizability of bias measurements across generation tasks. Further, we find little evidence that the relationship between bias and scale is stable across different evaluations, even in these evaluations using the same metrics applied to the same type of bias (i.e. gender-occupation).
5 Discussion and Limitations
This study highlights that the adaptability of LLMs to a wide range of downstream tasks—their core strength—also represents a fundamental challenge for evaluation. Evidently, potentially unwanted correlations also change across tasks in a way that makes measuring the potential for downstream harm difficult. Real-world usage and harms from LLMs are quickly evolving, and the field of algorithmic fairness will need to keep pace in order to contribute to the mitigation of real-world harm. In particular, we suggest moving away from the common “trick tests” used today and towards RUTEd evaluations that match real-world use cases and have articulable associations with the real-world harms we aim to address.
This is a limited study in many respects. We only calculate these three metrics for seven models, three tasks, and one genre of bias evaluation: gender and occupation. Even within this genre, our reliance on the WinoBias sentences restricts us to a gender binary, certain occupations, and correlations solely between gender and occupation rather than with other variables (e.g., a gender stereotype of high-performers and low-performers in a single occupation). In part this restricted setting is to make comparison easier with past work, but with the RUTE distinction more developed, we hope future work will broaden the testing in each of these respects.
An additional important limitation is that there is still room for improvement to meet the ideal of RUTE. In our case, though we have based our evaluations on realistic use cases, the phrasing of the prompts was not based on a systematic review of different types of prompts that could be used for completing each task. A more RUTEd evaluation would incorporate language from different real prompts used for each given task. More publicly available data on real usage is needed to accomplish this.
6 Impact statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Example Model Outputs
Table A1 shows example output for each of the content generation tasks. We provide an abbreviated version of the example bedtime stories, as we limited them to 1000 tokens, so most are quite long.
Table A1: Example output
Context 7b-chat Output 13b-chat Output 70b-chat Output

### v2

@@ 3.2 RUTEd evaluation tasks
We include “Once upon a time” because, in initial trials without it, the model would sometimes generate text that discusses bedtime stories rather than immediately generating a particular story. We used a maximum length of 1000 tokens because this would be around 60 to 90 seconds spoken aloud at a slow-to-medium pace.
User personas
An increasingly popular and influential use case for LLMs has been generating synthetic data that approximates human behavior, such as in psychology research (Crockett and Messeri, 2023; Dillion et al., 2023; Harding et al., 2023). In human-computer interaction, researchers have been exploring the interaction between LLMs as a data source, including groups of LLMs to simulate social media behavior (Hämäläinen et al., 2023; Park et al., 2022, 2023).
User research typically involves testing user personas. These are short character sketches of fictional people that include information such as a name, profile picture, hobbies, accessibility needs, beliefs, and demographics (age, gender, education level, etc.). Bias in this context could impact whose needs are considered when a product is designed, built, and deployed. It could also lead to real-world gaps in who is best served by new products and technology, such as in the medical context, in which longstanding gender and racial biases in medical research and product development that have contributed to health disparities (Hoffman et al., 2016; Murthy et al., 2004).
To generate the user personas, we use the prompt:
@@
For this task, we use a maximum length of 150 tokens to reflect the typical length of user personas. While this resulted in sufficient detail, it sometimes led to a mid-sentence cutoff. Because we knew from the task that the text should be focused on a single individual (i.e., with a consistent gender), we did not need completed text to extract the required information to calculate bias metrics.
ESL learning exercises
One of the most prevalent and widely discussed LLM uses is education; for example, the popular online learning platform Khan Academy has heavily invested in Khanmigo, a GPT-4-powered learning assistant, which OpenAI has emphasized in their marketing (OpenAI, 2023). An important use case amenable to bias evaluation is the generation of instructional content such as explanations, examples, and test questions.
We prompt the model to generate sentences that could be used in instruction to demonstrate the use of pronouns to a language learner in the category of English as a Second Language (ESL). In this case, a significant gender imbalance could have similar harms as those of bedtime stories for children or for adults who are getting to know the culture and social norms of English-speaking regions, as well as harm through language learning as students learn to express different concepts at different rates and learn to favor certain words in certain contexts. To generate the sentences, we input the prompt:
@@
For this task, we use a maximum length of 100 tokens, corresponding to typical usage.
@@ 3.3 Metrics
We develop three metrics based on the extant literature. For the decontextualized sentence-completion task, we directly extract the normalized probability that the next word in the sentence containing occupation $o$ was a “male” word, $p_{o}^{m}$, or a “female” word, $p_{o}^{f}$.
@@
For the long-form text generation tasks, we must statistically estimate probabilities. For the Llama-2, GPT-4, and Mixtral-8x7B models, we generate $n=30$ replicates per task and occupation; for the Flan-PaLM models, we generate $n=64$. Models were set to default temperature with no minimum token probability and with the aforementioned maximum tokens for each context. Then, for each occupation, $o$, we calculate the proportion of replicates for that gender in which the generated text was about males, $\hat{p}_{o}^{m}$, and females, $\hat{p}_{p}^{f}$. Those for which greater than half of the pronouns refer to males are categorized as “male” replicates; the others are categorized as “female.” Because each occupation has an associated gender-majority, we also calculate the proportion of replicates that were gender-stereotypical, $\hat{p}_{o}^{s}$, and gender anti-stereotypical, $\hat{p}_{o}^{a}$. Replicates with no such pronouns are dropped. For clarity, we define metrics with the hatless notation and plug in $\hat{p}$ when necessary.
Neutrality
@@
We define the neutrality metric as $m^{\text{neutrality}}=\frac{1}{O}\sum_{o}\left|p^{m}_{o}-p^{f}_{o}\right|$. This metric is the one originally used in the BigBench Gender Sensitivity-English task (Srivastava et al., 2022). Essentially, this measures a distance from parity. When applying this metric to the decontextualized sentence completion task, this metric is zero if the male words and female words have equal probability of coming next in the sentence. When applied to the RUTEd long-form text generation, this metric is zero if male and female replicates are equally likely to be generated.
Skew
@@
Rather than the absolute difference from parity, we define the skew metric as the average tendency of the model to return male words or replicates instead of female words or replicates. $m^{\text{skew}}=\frac{1}{O}\sum_{o}\left(p^{\text{m}}_{o}-p^{\text{f}}_{o}\right)$. If male words/replicates have a higher probability systematically across all considered occupations, this metric is positive. Conversely, if female words/replicates have a systematically higher probability, the metric is negative.
Stereotype
Some studies have measured the difference between stereotypical and anti-stereotypical token generation (e.g., de Vassimon Manela et al., 2021; Nadeem et al., 2020). We define the stereotype metric analogously to neutrality: $m^{\text{stereotype}}=\frac{1}{O}\sum_{o}\left(p_{o}^{s}-p_{o}^{a}\right)$. To create a standard benchmark evaluation, the sum ranges over occupations in the benchmark. Here, $O$ represents the number of occupations used in the task.
@@ 3.4 Statistical uncertainty
Because the probabilities in the decontextualized evaluations are directly observed, there is no statistical uncertainty (and therefore no error bars in Figure 1). For the RUTEd evaluations, we estimate the sampling variance of each estimated probability. To calculate this, we first note that one component used in each metrics is $d_{o}=\hat{p}_{o}^{m}-\hat{p}_{o}^{f}=\hat{p}_{o}^{m}-(1-\hat{p}_{o}^{m})$. For the RUTEd tasks, using a simple plug-in estimator of the sampling variance of $\hat{p}_{o}^{m}$, we get $d_{o}=2\hat{p}_{o}^{m}-1$ with sampling variance $\hat{\sigma}_{o}^{2}=4\frac{\hat{p}_{o}^{m}(1-\hat{p}_{o}^{m})}{n}$. We apply this to each of the three metrics in the following two sections—where the variance of skew and stereotype are equal, and therefore only one derivation is needed.
Neutrality
A rule of thumb for assuming approximate normality of $\hat{p}$ for a binomial distribution is that it requires at least ten positive and negative examples (Peizer and Pratt, 1968). In our case, we largely meet or surpass this standard, so for mathematical convenience, we proceed under the assumption of normality—specifically, that $d_{o}\sim N(\mu_{o}=2\hat{p}_{o}^{m}-1,\hat{\sigma}^{2}_{o})$. This implies that $|d_{o}|$ has a folded normal distribution with mean $\mu_{Y,o}=\hat{\sigma}_{o}{\sqrt{\tfrac{2}{\pi}}}\,e^{(-\mu_{o}^{2}/2\ \hat{\sigma_{o}}^{2})}+\mu_{o}\left(1-2\,\Phi(-{\tfrac{\mu_{o}}{\hat{\sigma}_{o}}})\right)$ and variance $\sigma_{Y,o}=\mu_{o}^{2}+\ \hat{\sigma}^{2}_{o}-\mu_{Y,o}^{2}$. This implies that the sampling variance of $m_{R}^{\text{neutrality}}$ (where $R$ denotes a RUTEd evaluation) is given by $\frac{1}{O^{2}}\sum_{o}\sigma_{Y,o}$.
Skew and stereotype
Calculating the sampling variance of $m_{R}^{\text{skew}}$ and $m_{R}^{\text{stereotype}}$ is derived from $\hat{\sigma}_{o}^{2}$ by averaging across independent approximate normal distributions. Therefore, in both cases, the sampling variance of the estimator is $\frac{1}{O^{2}}\sum\hat{\sigma_{o}}^{2}$.
3.5 Models
We generated content and calculated metrics for models from four different families: Llama-2, Flan-PaLM, GPT-4-0125-preview, and Mixtral-8x7B. For the Llama-2 and Flan-PaLM models that have base models (i.e., only pre-trained) and chat (i.e., pre-trained and instruction-tuned) versions, we used the chat versions to mimic consumer use. For Llama-2 and Flan-PaLM, we evaluated several model sizes: the Llama-2 7, 13, and 70 billion parameter models (Touvron et al., 2023), and for Flan-PaLM, we evaluate the extra-small (XS), small (S), medium (M), and large (L) models (Chung et al., 2022). For GPT-4 and Mixtral-8x7B, we did not have access to next-word probabilities from the model providers, so we could not run the decontextualized standard benchmarks for these models, and therefore these models only contribute to the between-RUTEd evaluations analysis.
@@ 4 Results
We present five subsections of results: correlations between standard benchmarks and RUTEd evaluations, correlations across RUTEd evaluations, and three robustness checks: disaggregation by occupation, mode collapse, and prompt variation.
4.1 Correlations between standard benchmarks and RUTEd evaluations
Neutrality Skew Stereotype
Bedtime -0.07 0.57 0.36
Personas -0.25 0.54 -0.36
ESL 0.18 -0.39 0.54
Table 1: Rank correlation between standard benchmarks and RUTEd evaluations for each metric.
Figure 1: Results of 102 bias evaluations for three sizes of Llama-2 (blue), four sizes of Flan-PaLM (orange), GPT-4-0125-preview (green), and Mixtral-8x7B (purple), each on three metrics (neutrality, skew, stereotype) as a decontextualized standard benchmark and across three contexts (Bedtime Stories, User Personas, ESL Learning Exercises). Error bars indicate 95% confidence intervals. The standard benchmarks (top row) fail to predict the results of the RUTEd evaluations (other rows).
For each of the three metrics, there is little correlation between the standard benchmarks and any of the RUTEd evaluations. This is summarized in Table 1, which shows Spearman’s rank correlations. The average of the nine correlations is 0.12 with minimum -0.39 and maximum 0.57. For none of the metrics or RUTEd evaluations are the correlations consistently positive. When correlation is negative, ranking models by the standard benchmark evaluation would result in an ordering that is inversely related to the ordering based on the RUTEd evaluation.
In Figure 1, we display all 102 quantities. Columns of the grid correspond to metric types (i.e., neutrality, skew, and stereotype), and rows correspond to contexts (i.e., decontextualized, Bedtime Stories, personas, and ESL).
There is little consistency in model performance for any of the three metrics, as indicated by the rank correlations. We can consider particular cases in which a decision-maker would use the standard benchmarks. Consider, for example, if one were using a standard benchmark to select the least biased of the three sizes of Llama-2 (blue). On each of the three neutrality metrics, the standard benchmark results (i.e., the top row) assert that the 13B model is the least biased. However, on the nine RUTEd evaluations, only three of them show the 13B model as the least biased—exactly as many as we would expect by random chance.
A decision-maker may instead be selecting across all the models. For neutrality, the least biased is still Llama-2 13B. For skew, the least biased is Flan-PaLM L; note that for skew and stereotype, the values can be negative, and if some are, then still the lowest score (i.e., most negative score) is considered the least biased. For stereotype, the least biased is Flan-PaLM XL. Among the nine RUTEd evaluations, none of them assert the same as the corresponding standard benchmark. If we selected models at random, we would be correct approximately one in seven times, as we are excluding GPT-4 and Mixtral for this comparison.
4.2 Correlations between RUTEd evaluations
While the previous section showed that the standard benchmarks fail to reliably predict any of the three RUTEd evaluations, it is also worth considering whether the RUTEd evaluations can predict each other. If this were the case, then one RUTEd evaluation could be used to establish the bias of models in a more general sense. We largely found that this was not the case in our study, but we first discuss one pattern that emerged from the data: consistency across models, though not across model sizes.
Contexts Correlation
Bedtime Personas 0.042
Bedtime ESL 0.057
Personas ESL 0.183
Table 2: Rank correlation between RUTEd evaluations.
The inconsistency is clearer in the three pairwise correlations averaged across metrics, shown in Table 2, which shows Spearman’s rank correlation between each pair of RUTEd evaluations, averaged over the three potential metrics of interest. While each correlation is positive, they are near zero. This suggests that, in these three contexts, selecting or ranking models based on one context would not be a reliable way to identify the least biased models for application to another context. This echoes arguments for context-specific fairness from perspectives such as statistical theory (Anthis and Veitch, 2023), inverse reinforcement learning (Blandin and Kash, 2024), and social computing (Madaio et al., 2022).
4.3 Disaggregation by occupation
While bias metrics are typically calculated across individual terms, such as occupations, it is possible that there is correlation between standard benchmarks and RUTEd evaluations among occupations even though there is no correlation in aggregate. In detailed examination of the Llama-2 models, we do not find this to be the case, with more detail and visualizations in Appendix B.
4.4 Mode collapse
Mode collapse, a phenomenon in which a generative model produces only very similar outputs (Salimans et al., 2016), could distort bias estimates if the same replicate—possibly with small variation in wording—is generated repeatedly. We analyzed the 10,800 replicates for Llama-2 models (3 models, 3 RUTEd evaluations, 40 occupations, and 30 replicates per group) by first finding the groups of 30 replicates with the same model, evaluation, and occupation that had the highest average cosine similarity amongst themselves, using all-MiniLM-L6-v2 for sentence embeddings. We then manually inspected the groups with the most similarity and a random sample of other groups. We find a variety of replicates, even within the groups with the most cross-replicate similarity, suggesting that our findings are not the result of mode collapse. Further work in this area could introduce additional prompt variations as well as variation in model temperature or other hyperparameters.
4.5 Prompt variation
Because LLM output often varies across differently worded prompts with similar meanings (Dominguez-Olmedo et al., 2024; Salinas and Morstatter, 2024), we tested the Llama-2 models across 10 standard benchmark prompt templates (e.g., {occupation} is for ”) and 30 RUTEd prompt templates (10 for each context, e.g., a bedtime story about a child or a young person). We find that variation in the resultant metrics was significantly higher within standard benchmark results than within each RUTEd context. Second, we calculated the correlation across occupations, varying use of the original template or the mean result across all 10 templates. As shown in Appendix C, we found that standard benchmarks continue to have low correlation with RUTEd evaluation results, suggesting that our primary results are robust.
5 Conclusion
Our findings suggest that standard benchmarks are not robust to a relatively simple extension to realistic long-form text generation tasks, raising concerns about their continued use. We build on prior work that shows intrinsic metrics are poor predictors of extrinsic metrics (Cao et al., 2022) by showing that even extrinsic metrics fail, in this case, to predict tasks more grounded in real-world use. The adaptability of LLMs to diverse downstream tasks—their core strength—is a fundamental challenge for evaluation. Moreover, we find insufficient evidence to conclude that our three RUTEd evaluations are reliable predictors of each other. As real-world harms from LLMs quickly increase and evolve, we suggest moving away from these “trick tests” and towards RUTEd evaluations in the context of application. It is possible that more general benchmarks can be devised, but until then, we suggest that bias evaluation should be context-specific. At least, practitioners should not count on standard benchmarks when they decide which LLM to apply in their real-world contexts.
6 Limitations
While the present work lays a foundation for comparing standard benchmarks to RUTEd evaluations, more expansive development and testing is needed. As shown in Figure 1, we conducted tens of thousands of LLM trials that resulted in 102 gender-occupation bias quantities (i.e., combinations of three metrics, four evaluations contexts including decontextualized, and nine models—leading to 108 quantities, though we were unable to calculate three metrics for two models in the decontextualized evaluations due to limitations of public APIs, resulting in 102).
However, in each of these areas, this work should be expanded: more sensitive attributes (e.g., race), more social attributes (e.g., job applicant quality), more metrics, more contexts, or more models. This work should contend with the social complexities of other domains of bias as well as limitations of extant datasets (Blodgett et al., 2021). Even within the genre of gender-occupation bias, we are restricted to a gender binary, certain occupations, and correlations rather than other gender-occupation associations (e.g., a gender stereotype of the high-performers and low-performers within a single occupation). Examining new genres of bias could be more informative, but our goal was to show that there is instability in even this simple generalization from standard benchmarks to common LLM usage. This restricted setting allows us to make a targeted and cohesive argument based on the current literature, but it is limited in terms of the development of specific bias metrics that we would encourage be practically applied. We hope future work will further develop the RUTEd paradigm, such as taxonomizing the different dimensions in which an evaluation can be increasingly realistic. An example of a genre in which similar testing could be done is the association between race-associated names and employee performance. This domain has been less common in NLP than gender-occupation but has been a primary interest of economists in audit studies of employer bias (e.g., Bertrand and Mullainathan, 2004; Veldanda et al., 2023).
An important limitation is that, though it was important to show that decontextualized evaluations fail to correlate with even an analogous long-form text generation, there is still room for improvement to meet the ideal of RUTEd evaluations. In our case, though we have based our evaluations on realistic use cases and have posited tangible effects that could occur, we did not conduct tests with the widely varied prompts (e.g., syntax, language, additional information) that are present in real-world LLM use. It will be particularly important to consider datasets of real-world interactions, such as WildChat (Zhao et al., 2024), when constructing such evaluations. It will also be important to test for tangible effects, though the empirical demands of such research will be significant.
Finally, we note that while our results suggest caution when using standard bias benchmarks in real-world application, they do not diminish the contributions or usefulness of these benchmarks or other prior work. The field of algorithmic fairness has built technical and empirical frameworks step-by-step, and shifting paradigms (e.g., towards decoder-only transformer architectures) have made it challenging to keep pace with real-world use.
Appendix A Disaggregation by occupation
While our focus is the aggregate measure across the 40 tested occupations, we also examined the results across occupations based on the numerous trials conducted for each. Like in the aggregate, there is little correlation between standard benchmarks and RUTEd evaluations. We show the disaggregation for skew in Figure A1, alongside analogous figures for stereotype and neutrality. We find that both methods reveal similar occupations with highly female-skewed output (e.g., housekeeper, receptionist) and with highly male-skewed output (e.g., construction worker, carpenter). However, based on the RUTEd evaluations, we find that standard benchmarks overestimate the relative skew of the most female-skewed occupations and view some of the middling occupations (i.e., not the most female- or male-skewed), such as lawyer and baker, as relatively more male-skewed. We include a walkthrough of the figure; again, the patterns observed in the disaggregate analysis were not our focus, but this may be an important approach for future research and RUTEd evaluations that focus on particular occupations, such as for fairness in an occupation-specific LLM application.
We show results disaggregated across the 40 occupations for the three Llama-2 models in Figure A1, Figure A2, and Figure A3. For clarity, we briefly walk through the skew figure and the pattern of relative skew across occupations (Figure A1.
  1. 1.
    First, notice that for comparability between standard benchmarks and RUTEd evaluations across occupations, these figures, but not the main text, report normalized metrics (i.e., $\mu=0,\sigma=1$). The quantities in the figures are not directly comparable to those in the main text.
  2. 2.
    Second, notice that the occupations are ordered from the most male-skewed at the top to the most female-skewed at the bottom, which is reflected in the positions of the purple and gray marks in the scatterplot.
  3. 3.
    Third, notice that the horizontal bars reflect the difference between skew as measured by the standard benchmarks and that as measured by the RUTEd evaluations. For the most male-skewed occupation, construction worker, it was among the most male-skewed for both standard benchmarks and RUTEd evaluations. The coral-colored bar of that row indicates a negative difference. In other words, the RUTEd evaluations show this is relatively male-skewed compared to what the standard benchmarks indicate.
  4. 4.
    Fourth, notice that the largest coral bars are near the bottom of the y-axis, and the largest blue bars are near the middle of the y-axis. In other words, if we only had standard benchmarks, we would judge that the models tend to skew even further towards female for the most female-skewed occupations (e.g., housekeeper, receptionist), which, informally, seem to be stereotypically associated with female gender. We would also judge that, on average, the models tend to skew towards male not for the most male-skewed occupations (again, informally, this would be construction worker, carpenter, etc.) but for middling or moderately male-skewed occupations. This is only a speculative, exploratory analysis, but we encourage future work that disaggregates across occupations.
Figure A1: Skew metrics disaggregated by occupation for the three Llama-2 models. On the left, the bar chart shows the normalized difference between the average of standard benchmark skew evaluations and the average of RUTEd skew evaluations. The difference is displayed as a number next to the occupation as well as the magnitude of the bar, and the occupations are ordered by the average between skew across the standard benchmarks and skew across the RUTEd evaluations (both equally weighted). On the right, the scatterplot shows the exact skew values for 12 evaluations per occupation (3 models, 4 contexts). Shapes correspond to different sizes of the Llama-2 model. The standard benchmarks are shown in purple. All RUTEd evaluations are shown in gray.
Figure A2: Stereotype metrics disaggregated by occupation for the three Llama-2 models. On the left, the bar chart shows the normalized difference between the average of standard benchmark stereotype evaluations and the average of RUTEd stereotype evaluations. The difference is displayed as a number next to the occupation as well as the magnitude of the bar, and the occupations are ordered by the average between stereotype across the standard benchmarks and sterotype across the RUTEd evaluations (both equally weighted). On the right, the scatterplot shows the exact stereotype values for 12 evaluations per occupation (3 models, 4 contexts). Shapes correspond to different sizes of the Llama-2 model. The standard benchmarks are shown in purple. All RUTEd evaluations are shown in gray.
Figure A3: Neutrality bias metrics disaggregated by occupation for the three Llama-2 models. On the left, the bar chart shows the normalized difference between the average of standard benchmark neutrality evaluations and the average of RUTEd neutrality evaluations. The difference is displayed as a number next to the occupation as well as in the magnitude of the bar, and the occupations are ordered by the average between neutrality bias across the standard benchmarks and neutrality bias across the RUTEd evaluations (both equally weighted). On the right, the scatterplot shows the exact neutrality bias values for 12 evaluations per occupation (3 models, 4 contexts). Shapes correspond to different sizes of the Llama-2 model. The standard benchmarks are shown in purple. All RUTEd evaluations are shown in gray.
Appendix B Prompt variation
Table B1, Table B2, and Table B3 show the templates used for prompt variation tests. Figure B1 shows the correlations between templates within each context (No Context, Bedtime, Personas, ESL). Table B4 shows 108 rank correlations between standard and RUTEd for the 40 occupations in each combination of context, metric, template combination, and model for the Llama-2 models. Correlations across occupations tend to be stronger than those correlations across models, but the correlations still tend to be close to zero for predicting the rank order of occupations of RUTEd results from standard results.